# Dynamical Dark Energy Beyond Planck? Constraints from multiple CMB probes, DESI BAO and Type-Ia Supernovae

### v1

@@ Abstract
Baryon Acoustic Oscillation (BAO) measurements from the Dark Energy Spectroscopic Instrument (DESI) collaboration, when combined with Planck satellite Cosmic Microwave Background (CMB) data and Type Ia Supernovae, suggest a preference for Dynamical Dark Energy (DDE) at a significance level ranging from $2.5\sigma$ to $3.9\sigma$. In this work, I test whether, and to what extent, this preference is supported by CMB experiments other than Planck. I analyze the Atacama Cosmology Telescope (ACT) and South Pole Telescope (SPT) temperature, polarization, and lensing spectra at small scales, eventually combining them with Planck or WMAP 9-year observations at large angular scales. My analysis shows that ACT and WMAP data, when combined with DESI BAO and Pantheon-plus Supernovae, yield independent constraints with a precision comparable to Planck. Notably, in this case, the cosmological constant value is recovered within two standard deviations. A preference for DDE reappears when Pantheon-plus is replaced with distance moduli measurements from the Dark Energy Survey Supernova program (DESy5). However, it remains less pronounced compared to the Planck-based results. When considering SPT data, no clear preference for DDE is found, although the parameter uncertainties are significantly larger compared to both Planck- and ACT-based constraints. Overall, CMB experiments other than Planck generally weaken the evidence for DDE. I argue that the subsets of Planck data that strengthen the shift toward DDE are the temperature and E-mode polarization anisotropy measurements at large angular scales $\ell\lesssim 30$.
@@ I Introduction
The Dark Energy Spectroscopic Instrument (DESI) has recently released measurements of Baryon Acoustic Oscillations (BAO) based on its first year of observations, which include data from over 6 million extra-galactic objects, such as galaxies, quasars, and Lyman-$\alpha$ forest tracers Adame et al. (2024a, b); Lodha et al. (2024); Adame et al. (2024c). DESI BAO provide precise constraints on the transverse comoving distance, the Hubble rate, and their combination (all relative to the value of the sound horizon at the drag epoch) across seven redshift bins spanning a range $4.20.1<z<$ 4.2.
@@
One of the most important – yet preliminary – results emerging from these new observations concerns the nature of Dark Energy (DE). [^1] Other potential momentous implications concern neutrino cosmology. The upper limit on the total neutrino mass resulting from DESI is approaching the lower limit imposed by oscillation experiments. For discussions on the implications for neutrinos, including their total mass and hierarchy, we refer to Refs. Adame et al. (2024a); Craig et al. (2024); Wang et al. (2024a); Allali and Notari (2024); Sen and Smirnov (2024); Green and Meyers (2024); Elbers et al. (2024); Naredo-Tuero et al. (2024); Jiang et al. (2024a); Rebouças et al. (2024). As initially emphasized by the DESI collaboration and corroborated by multiple independent re-analyses Adame et al. (2024a); Lodha et al. (2024), combining DESI BAO with Planck Cosmic Microwave Background (CMB) data and Type Ia Supernovae (SN) distance moduli measurements – from either the Pantheon-plus catalog Scolnic et al. (2022); Brout et al. (2022), the Union3 compilation Rubin et al. (2023), or the five-year observations of the Dark Energy Survey (DESy5) Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024) – yields moderate to strong evidence for a time-evolving DE component, often referred to as Dynamical Dark Energy (DDE). Using the well-known Chevallier-Polarski-Linder (CPL) parameterization Chevallier and Polarski (2001); Linder (2003), where the DE equation of state (EoS) is given by $w(a)=w_{0}+w_{a}(1-a)$, the data show a preference for a present-day quintessence-like EoS ($w_{0}>-1$) that crossed the phantom barrier ($w_{a}<0$). This preference reaches a significance level of about $3.9\sigma$ when combining Planck CMB, DESI BAO and DESy5 SN, while it is reduced to around $2.5\sigma$ when replacing DESy5 with Pantheon-plus Adame et al. (2024a).
@@
If confirmed, this result would represent the first concrete evidence of new physics beyond the standard model of cosmology, potentially bringing a significant portion of the theoretical high-energy physics community back to the board. Therefore, somewhat unsurprisingly, following the DESI data release, a multitude of reanalyses and attempts at physical interpretation of this signal have heated up the debate Cortês and Liddle (2024); Patel and Amendola (2024); Orchard and Cárdenas (2024); Liu et al. (2024); Chudaykin and Kunz (2024); Notari et al. (2024); Gialamas et al. (2024); Wang et al. (2024b); Lodha et al. (2024); Wang and Piao (2024); Giarè et al. (2024a); Carloni et al. (2024); Colgáin et al. (2024); Tada and Terada (2024); Yin (2024); Luongo and Muccino (2024); Park et al. (2024); Wang et al. (2024c); Shlivko and Steinhardt (2024); Ye et al. (2024); Li et al. (2024); Yang et al. (2024); Jiang et al. (2024b); Giarè et al. (2024b); Poulin et al. (2024); Pourojaghi et al. (2024a); Dinda and Maartens (2024); Escamilla et al. (2024a); Sabogal et al. (2024); Ghedini et al. (2024); Wolf et al. (2024); Arjona and Nesseris (2024). Given its potential impact on our understanding of the Universe, caution is certainly mandatory and thoroughly testing the robustness of these findings is of paramount importance.
@@
A crucial aspect that has already undergone significant cross-checking is the role played by the parameterization used to describe the DE EoS Hernández-Almada et al. (2024); Pourojaghi et al. (2024b); Ramadan et al. (2024); Berghaus et al. (2024); Qu et al. (2024a); Notari et al. (2024); Adolf et al. (2024). In Ref. Giarè et al. (2024b), we noted that assuming the CPL parameterization is not the primary driver of this preference. When combining CMB, DESI BAO, and SN measurements, the same preference for a present-day quintessence-like DDE that crosses the phantom barrier consistently emerges across a wide range of different parameterizations. In this sense, the preference is robust against different models.
@@
With the parameterization issue clarified, the next step is to investigate potential observational systematic effects in the datasets involved. Systematic effects in DESI BAO measurements have been a topic of debate. Notably, the DESI BAO measurement at $z=0.71$ shows approximately $3\sigma$ tension with predictions based on the Planck best-fit $\Lambda$CDM cosmology. This outlier could potentially be responsible for part of the DESI signals for new physics, including (part of) the preference for DDE Wang et al. (2024d); Colgáin et al. (2024); Naredo-Tuero et al. (2024).
@@
Similarly, the impact of SN measurements was recently highlighted in Ref. Efstathiou (2024), where cross-correlating the Pantheon-plus and DESy5 supernova samples suggested evidence for a calibration difference of $\sim 0.04$ mag between low and high redshifts. Correcting for this offset brings the DESy5 sample into closer agreement with Planck’s $\Lambda$CDM cosmology. Since the parameter range favored by the uncorrected DESy5 sample diverges from many other cosmological datasets, it has been suggested that the evidence for DDE might primarily be due to (systematics in) DESy5 SN.
@@
To the best of my knowledge, so far all the analyses surrounding the DESI and SN preference for DDE have primarily (although somewhat unconsciously) relied on temperature and polarization anisotropy spectra measured by the Planck satellite Aghanim et al. (2020a, b), as well as lensing power spectrum reconstructions from Planck, whether from Plik Aghanim et al. (2020c) or the more recent NPIPE-PR4 Carron et al. (2022) likelihoods. [^2] See recent Ref. Roy Choudhury and Okumura (2024) for updated constraints derived from Planck-PR4 based likelihoods. The only notable non-Planck contribution to the CMB side of these analyses comes from the Atacama Cosmology Telescope (ACT) lensing likelihood (ACT-DR6Madhavacheril et al. (2024); Qu et al. (2024b), often used in conjunction with NPIPE-PR4 or Plik lensing. However, this seems a modest usage of the multitude of CMB observations available beyond Planck. Ground-based CMB experiments, such as ACT Aiola et al. (2020); Madhavacheril et al. (2024); Qu et al. (2024b) and the South Pole Telescope (SPT) Dutcher et al. (2021); Balkenhol et al. (2023), provide measurements of temperature and polarization anisotropies that, while partially overlapping with Planck, extend into higher multipole regions (i.e., smaller angular scales). When considered independently and especially in combination with other large-scale CMB data (such as from WMAP Hinshaw et al. (2013)), these experiments can achieve a level of precision on cosmological parameters approaching that of Planck.
@@
Expanding the analysis and testing whether the preference for DDE is confirmed by CMB measurements other than Planck is certainly feasible and worthwhile for a good number of well-motivated reasons. Firstly, over the years, several mild anomalies have been identified in the Planck CMB measurements, many of which affect parameters important for the late-time dynamics of the Universe. One notable example concerns the higher lensing amplitude inferred from Planck temperature and polarization spectra, captured by the phenomenological parameter $A_{\rm L}$ Calabrese et al. (2008), which deviates from the baseline value ($A_{\rm L}=1$) by about $2.8\sigma$ Aghanim et al. (2020a); Di Valentino et al. (2016); Renzi et al. (2018); Domènech et al. (2020). This issue can be recast into a well-documented preference for a closed Universe, as discussed in several recent works Park and Ratra (2019); Handley (2021); Di Valentino et al. (2019); Efstathiou and Gratton (2020); Di Valentino et al. (2021a); Benisty and Staicova (2021); Vagnozzi et al. (2021a, b); Di Valentino et al. (2021b); Yang et al. (2021); Cao et al. (2021); Dhawan et al. (2021); Dinda (2022); Gonzalez et al. (2021); Akarsu et al. (2023); Cao and Ratra (2022); Glanville et al. (2022); Bel et al. (2022); Yang et al. (2023); Stevens et al. (2023); Favale et al. (2023). Secondly, focusing solely on Planck CMB data, one might speculate about a latent preference for a phantom-like DE component Aghanim et al. (2020a); Escamilla et al. (2024b) that is closely related to large-scale temperature and E-modes polarization measurements Escamilla et al. (2024b); Giarè et al. (2024c); Ben-Dayan et al. (2024). This latent preference (not confirmed by other CMB experiments) was previously washed out by older BAO measurements from the SDSS surveys Alam et al. (2021). However, it is worth checking whether it now sums up with the DESI and SN tendency for a phantom DDE, thereby strengthening its preference. Last but not least, discrepancies often arise between Planck and Planck-independent experiments as reported by different groups Lin and Ishak (2021); Forconi et al. (2021); Handley and Lemos (2021); La Posta et al. (2023); Di Valentino et al. (2023, 2022); Giarè et al. (2023a); Calderón et al. (2023); Giarè (2023). Most disagreements involve small-scale CMB measurements from Data Release 4 of the ACT (ACT-DR4), and concern the value of inflationary parameters (most notably the value of the spectral index Giarè et al. (2023a, b)). However, differences of $\sim 1.5\sigma$ to $2\sigma$ can potentially propagate to other cosmic parameters, such as the matter density parameter and the sound horizon Giarè (2023), which could have implications for DDE that warrant further investigation.
@@
Driven by these motivations and concerns, I present the results of a comprehensive analysis aimed at clarifying the role of various CMB experiments in conjunction with DESI BAO and SN measurements. My analysis shows that the preference for DDE is reduced when considering CMB experiments beyond Planck. Specifically, when ACT data (alone or in combination with WMAP) are analyzed together with DESI BAO and Pantheon-plus, there is no significant preference for DDE, and the cosmological constant model falls within the 95% confidence level (CL) contours. On the other hand, a preference for DDE is still observed when combining (WMAP+)ACT with DESI BAO and DESy5, confirming the importance of the latter SN catalouge. However, the statistical significance of this preference is diminished compared to Planck-based results. When considering SPT data, I find no convincing preference for DDE. In this case, the cosmological constant model consistently falls within the 95% CL contours, although parameter uncertainties are notably larger compared to those from Planck and ACT. I also contend that the temperature and E-mode polarization anisotropy measurements at large angular scales ($\ell\lesssim 30$) in the Planck data are the subset mostly responsible for reinforcing the shift towards DDE.

### v2

@@ Abstract
The latest Baryon Acoustic Oscillation (BAO) measurements from the Dark Energy Spectroscopic Instrument (DESI) collaboration, when combined with Planck satellite Cosmic Microwave Background (CMB) data and Type Ia Supernovae, suggest a preference for Dynamical Dark Energy (DDE) at a significance level ranging from $2.8\sigma$ to $4.2\sigma$. In this work, I test whether, and to what extent, this preference is supported by CMB experiments other than Planck. I analyze the latest Atacama Cosmology Telescope (ACT) and South Pole Telescope (SPT) temperature, polarization, and lensing spectra at small scales, eventually combining them with Planck or WMAP 9-year observations at large angular scales. My analysis shows that ACT and WMAP data, when combined with DESI BAO and Pantheon-plus Supernovae, yield independent constraints with a precision comparable to Planck. Notably, in this case, the cosmological constant value is recovered within two standard deviations. A preference for DDE reappears when Pantheon-plus is replaced with distance moduli measurements from the Dark Energy Survey Supernova program (DESy5). However, it remains less pronounced compared to the Planck-based results. When considering SPT data, no clear preference for DDE is found in combinations involving Pantheon-plus Supernovae, and the preference is significantly weaker in combinations involving DESy5. Overall, CMB experiments other than Planck generally weaken the evidence for DDE. I argue that the subsets of Planck data that strengthen the shift toward DDE are the temperature and E-mode polarization anisotropy measurements at large angular scales $\ell\lesssim 30$.
@@ I Introduction
The Dark Energy Spectroscopic Instrument (DESI) has recently released measurements of Baryon Acoustic Oscillations (BAO) based on its second year of observations (DR2), which include data from over 14 million extra-galactic objects, such as galaxies, quasars, and Lyman-$\alpha$ forest tracers Adame et al. (2025a, b); Lodha et al. (2025a); Adame et al. (2025c); Calderon et al. (2024); Abdul Karim et al. (2025a, b); Andrade et al. (2025). DESI BAO provide precise constraints on the transverse comoving distance, the Hubble rate, and their combination (all relative to the value of the sound horizon at the drag epoch) across seven redshift bins spanning a range $4.20.1<z<$ 4.2.
@@
One of the most important – yet preliminary – results emerging from these observations concerns the nature of Dark Energy (DE). [^1] Other potential momentous implications concern neutrino cosmology. The upper limit on the total neutrino mass resulting from DESI is approaching the lower limit imposed by oscillation experiments. For discussions on the implications for neutrinos, including their total mass and hierarchy, we refer to Refs. Adame et al. (2025a); Craig et al. (2024); Wang et al. (2024a); Allali and Notari (2024); Sen and Smirnov (2025); Green and Meyers (2025); Elbers et al. (2025); Naredo-Tuero et al. (2024); Jiang et al. (2025); Rebouças et al. (2025); Capozzi et al. (2025); Abdul Karim et al. (2025a). As initially emphasized by the DESI collaboration and corroborated by multiple independent re-analyses Adame et al. (2025a); Lodha et al. (2025a); Abdul Karim et al. (2025a); Lodha et al. (2025b), combining DESI BAO with Planck Cosmic Microwave Background (CMB) data and Type Ia Supernovae (SN) distance moduli measurements – from either the Pantheon-plus catalog Scolnic et al. (2022); Brout et al. (2022), the Union3 compilation Rubin et al. (2023), or the five-year observations of the Dark Energy Survey (DESy5) Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024) – yields moderate to strong evidence for a time-evolving DE component, often referred to as Dynamical Dark Energy (DDE). Using the well-known Chevallier-Polarski-Linder (CPL) parameterization Chevallier and Polarski (2001); Linder (2003), where the DE equation of state (EoS) is given by $w(a)=w_{0}+w_{a}(1-a)$, the data show a preference for a present-day quintessence-like EoS ($w_{0}>-1$) that crossed the phantom barrier ($w_{a}<0$). This preference reaches a significance level of about $4.2\sigma$ when combining Planck CMB, DESI BAO and DESy5 SN, while it is reduced to around $2.8\sigma$ when replacing DESy5 with Pantheon-plus Adame et al. (2025a).
@@
If confirmed, this result would represent the first concrete evidence of new physics beyond the standard model of cosmology, potentially bringing a significant portion of the theoretical high-energy physics community back to the board. Therefore, somewhat unsurprisingly, following the DESI data release, a multitude of reanalyses and attempts at physical interpretation of this signal have heated up the debate, see, e.g., Refs. Cortês and Liddle (2024); Patel et al. (2024); Orchard and Cárdenas (2024); Liu et al. (2024); Chudaykin and Kunz (2024); Notari et al. (2024); Gialamas et al. (2025); Wang et al. (2025); Lodha et al. (2025a); Wang and Piao (2024); Giarè et al. (2024a); Carloni et al. (2025); Colgáin et al. (2024a); Bhattacharya et al. (2024); Tada and Terada (2024); Yin (2024); Luongo and Muccino (2024); Park et al. (2024); Wang et al. (2024b); Shlivko and Steinhardt (2024); Ye et al. (2025); Li et al. (2024); Yang et al. (2024); Jiang et al. (2024); Giarè et al. (2024b); Poulin et al. (2025); Pourojaghi et al. (2025); Dinda and Maartens (2025); Escamilla et al. (2024a); Sabogal et al. (2024); Ghedini et al. (2024); Wolf et al. (2024); Arjona and Nesseris (2024); Wolf et al. (2025a); Chakraborty et al. (2025, 2025); Silva et al. (2025); Pan et al. (2025); Pan and Ye (2025); Ormondroyd et al. (2025); Wolf et al. (2025b); Anchordoqui et al. (2025); Pang et al. (2025); Colgáin et al. (2025a); Kessler et al. (2025); Nesseris et al. (2025); Shah et al. (2025); Santos et al. (2025); Brandenberger (2025); Bansal and Huterer (2025); Lewis and Chamberlain (2025); Roy Choudhury (2025); Teixeira et al. (2025). Given its potential impact on our understanding of the Universe, caution is certainly mandatory and thoroughly testing the robustness of these findings is of paramount importance.
@@
A crucial aspect that has already undergone significant cross-checking is the role played by the parameterization used to describe the DE EoS Calderon et al. (2024); Hernández-Almada et al. (2024); Malekjani et al. (2025); Ramadan et al. (2024); Berghaus et al. (2024); Qu et al. (2025); Notari et al. (2024); Adolf et al. (2024). In Ref. Giarè et al. (2024b), we noted that assuming the CPL parameterization is not the primary driver of this preference. When combining CMB, DESI BAO, and SN measurements, the same preference for a present-day quintessence-like DDE that crosses the phantom barrier consistently emerges across a wide range of different parameterizations. In this sense, the preference is robust against different models, as subsequently confirmed by the DESI collaboration itself Lodha et al. (2025b).
@@
With the parameterization issue clarified, the next step is to investigate potential observational systematic effects in the datasets involved. Systematic effects in DESI BAO measurements have been a topic of debate. Notably, the DESI BAO measurement at $z=0.71$ shows approximately $3\sigma$ tension with predictions based on the Planck best-fit $\Lambda$CDM cosmology. This outlier is potentially responsible for part of the DESI signals for new physics, including (part of) the preference for DDE Wang et al. (2024c); Colgáin et al. (2024a); Naredo-Tuero et al. (2024).
@@
Similarly, the impact of SN measurements was recently highlighted in Ref. Efstathiou (2025), where cross-correlating the Pantheon-plus and DESy5 supernova samples suggested evidence for a calibration difference of $\sim 0.04$ mag between low and high redshifts. [^2] See also Refs. Colgáin et al. (2022, 2024b); Malekjani et al. (2024) for discussions on the consistency of SN measurements across high and low redshifts, as well as Ref. Colgáin et al. (2025b) for other discussions surrounding potential calibration issues in DESy5. for other discussions surrounding Correcting for this offset brings the DESy5 sample into closer agreement with Planck’s $\Lambda$CDM cosmology. Since the parameter range favored by the uncorrected DESy5 sample diverges from many other cosmological datasets, it has been suggested that the evidence for DDE might primarily be due to (systematics in) DESy5 SN Huang et al. (2024); Efstathiou (2025); Notari et al. (2025). In response to these concerns, however, the DES collaboration in Ref. Vincenzi et al. (2025) showed that the debated $\sim 0.04$ mag offset between Pantheon-plus and DESy5 at low and high redshift is partly attributable to improvements in the modeling of supernova intrinsic scatter and host galaxy properties in DESy5 (accounting for up to $\sim 43\%$ of the offset), and partly ($\sim 38\%$) to a misleading comparison. The latter stems from the fact that different selection functions characterize the DES subsets included in Pantheon-plus and DESy5, leading to differences in individual supernova distance measurements due to distinct bias corrections.
@@
To the best of my knowledge, so far all the analyses surrounding the DESI and SN preference for DDE have primarily (although somewhat unconsciously) relied on temperature and polarization anisotropy spectra measured by the Planck satellite Aghanim et al. (2020a, b), as well as lensing power spectrum reconstructions from Planck, whether from Plik Aghanim et al. (2020c) or the more recent NPIPE-PR4 Carron et al. (2022) likelihoods. [^3] See recent Ref. Roy Choudhury and Okumura (2024) for updated constraints derived from Planck-PR4 based likelihoods. The only notable non-Planck contribution to the CMB side of these analyses comes from the Atacama Cosmology Telescope (ACT) lensing likelihood (ACT-DR6Madhavacheril et al. (2024); Qu et al. (2024), often used in conjunction with NPIPE-PR4 or Plik lensing. However, this seems a modest usage of the multitude of CMB observations available beyond Planck. Ground-based CMB experiments, such as ACT Aiola et al. (2020); Madhavacheril et al. (2024); Qu et al. (2024); Louis et al. (2025) and the South Pole Telescope (SPT) Dutcher et al. (2021); Balkenhol et al. (2023), provide measurements of temperature and polarization anisotropies that, while partially overlapping with Planck, extend into higher multipole regions (i.e., smaller angular scales). When considered independently and especially in combination with other large-scale CMB data (such as from WMAP Hinshaw et al. (2013)), these experiments can achieve a level of precision on cosmological parameters approaching that of Planck.
@@
Expanding the analysis and testing whether the preference for DDE is confirmed by CMB measurements other than Planck is certainly feasible and worthwhile for a good number of well-motivated reasons. Firstly, over the years, several mild anomalies have been identified in the Planck CMB measurements, many of which affect parameters important for the late-time dynamics of the Universe. One notable example concerns the higher lensing amplitude inferred from Planck temperature and polarization spectra, captured by the phenomenological parameter $A_{\rm L}$ Calabrese et al. (2008), which deviates from the baseline value ($A_{\rm L}=1$) by about $2.8\sigma$ Aghanim et al. (2020a); Di Valentino et al. (2016); Renzi et al. (2018); Domènech et al. (2020). This issue can be recast into a well-documented preference for a closed Universe, as discussed in several recent works Park and Ratra (2019); Handley (2021); Di Valentino et al. (2019); Efstathiou and Gratton (2020); Di Valentino et al. (2021a); Benisty and Staicova (2021); Vagnozzi et al. (2021a, b); Di Valentino et al. (2021b); Yang et al. (2021); Cao et al. (2021); Dhawan et al. (2021); Dinda (2022); Gonzalez et al. (2021); Akarsu et al. (2023); Cao and Ratra (2022); Glanville et al. (2022); Bel et al. (2022); Yang et al. (2023); Stevens et al. (2023); Favale et al. (2023). Secondly, focusing solely on Planck CMB data, one might speculate about a latent preference for a phantom-like DE component Aghanim et al. (2020a); Escamilla et al. (2024b) that is closely related to large-scale temperature and E-modes polarization measurements Escamilla et al. (2024b); Giarè et al. (2024c); Ben-Dayan et al. (2025). This latent preference (not confirmed by other CMB experiments) was previously washed out by older BAO measurements from the SDSS surveys Alam et al. (2021). However, it is worth checking whether it now sums up with the DESI and SN tendency for a phantom DDE, thereby strengthening its preference. Last but not least, discrepancies often arise between Planck and Planck-independent experiments as reported by different groups Lin and Ishak (2021); Forconi et al. (2021); Handley and Lemos (2021); La Posta et al. (2023); Di Valentino et al. (2023, 2022); Giarè et al. (2023a); Calderón et al. (2023); Giarè (2023); Giarè et al. (2024d); Gariazzo et al. (2025); Forconi and DI Valentino (2025). Most disagreements involve small-scale CMB measurements from Data Release 4 of the ACT (ACT-DR4), and concern the value of inflationary parameters (most notably the value of the spectral index Giarè et al. (2023a, b); Forconi and DI Valentino (2025)). While these problems have been largely mitigated in the latest Data Release 6 of the ACT temperature and polarization spectra (ACT-DR6), differences are still observed and can potentially propagate to other cosmological parameters, such as the matter density parameter and the sound horizon Giarè (2023), which could have implications for DDE that warrant further investigation.
@@
Driven by these motivations and concerns, I present the results of a comprehensive analysis aimed at clarifying the role of various latest CMB experiments in conjunction with DESI-DR2 BAO and SN measurements. My analysis shows that the preference for DDE is reduced when considering CMB experiments beyond Planck. Specifically, when the latest ACT-DR6 data (alone or in combination with WMAP) are analyzed together with DESI-DR2 BAO and Pantheon-plus, there is no strong preference for DDE, and the cosmological constant model essentially falls within or close to the 95 % confidence level (CL) contours. On the other hand, a preference for DDE is still observed when combining (WMAP+)ACT-DR6 with DESI-DR2 BAO and DESy5, confirming the importance of the latter SN catalogue. However, the statistical significance of this preference is diminished compared to Planck-based results. Similarly, when considering SPT data, I find no convincing preference for DDE in combinations that involve Pantheon-plus, and the preference becomes modest in combinations that involve DESy5. I also contend that the temperature and E-mode polarization anisotropy measurements at large angular scales ($\ell\lesssim 30$) in the Planck data are the subset mostly responsible for reinforcing the shift towards DDE.